OpenStudy (anonymous):

How many grams of magnesium metal will react completely with 6.3 liters of 4.5 M HCl? Show all of the work needed to solve this problem. Mg (s) + 2 HCl (aq) MgCl2 (aq) + H2 (g)

OpenStudy (chmvijay):

Mg + 2HCl -----> MgCl2 + H2 M = mol / L mol = M x L = 4.5 mol/L x 6.3 L = 28.4 mol HCl 1 mol Mg reacts with 2 mol HCl 28.4 mol HCl x (1 mol Mg / 2 mol HCl) = 14.2 mol Mg Molar mass of Mg is 24.3 g/mol. 14.2 mol Mg x (24.3 g / 1 mol) = 344 g Mg